A 70 turn, 10.0 cm diameter coil rotates at an angular velocity of 8.00 rad/s in a 1.35 T field, starting with the plane of the coil parallel to the field. Assume that the positive max emf is reached first.

(a) What is the peak emf?

=____ V

(b) At what time is the peak emf first reached?

=____ s

(c) At what time is the emf first at its most negative?

=____ s

(d) What is the period of the AC voltage output?

=____ s
Answer #1

E = BANωsinθ = BANωsin(ωt)

peak when sin(ωt) = 1

peak emf =  BANω

a) peak emf = BANω = 1.35*pi( 0.1/2)^2*70*8 = 5.93 V

b) T = 2pi /w = 2pi / 8 = 0.785398163 sec

first time peak = T /4 = 0.785398163 /4 =0.196349541 sec

c) T/2 =0.392699081

d ) period =T = 0.785398163 sec